Suppose Komla throws a die repeatedly until he gets a six. What is the probability
that he needs to throw more than 10 times to get a six, to 3 decimal places?
1
Expert's answer
2020-06-02T18:49:26-0400


Let X = the random variable denoting the number of throws Komla needs to get a six


Since, the probability of getting a six when a die is thrown is 1/6, then we have,


X ~ Geometric(p = 1/6)


The p.m.f. of X is given by,


P(X = x) = "\begin{cases}
 (1 - p)^{x-1}p, &\text{for }x = 1, 2, 3, .... \\
 0, &\text{otherwise}
\end{cases}"


"\therefore" The probability that Komla needs to throw more than 10 times to get a six


= P(X > 10)


= 1 - P(X "\leq" 10) [since the total probability is 1]


= 1 - "\displaystyle\sum_{x=1}^{10}P(X=x)"

= 1 - "\displaystyle\sum_{x=1}^{10}(1 - p)^{x-1}p"

= 1 - "p[1 + (1 - p) + (1 - p)^2 + ... + (1 - p)^9]"


= 1 - "p\frac{1-(1-p)^{10}}{1-(1-p)}"


= 1 - "p\frac{1-(1-p)^{10}}{p}"


= 1 - 1 + (1 - p)10


= (1 - p)10 = (1 - 1/6)10 = (5/6)10 = 0.162 (rounded to 2 decimal places)


Answer: The probability that Komla needs to throw more than 10 times to get a six is 0.162.